PRESSURE AND TEMPERATURE DEPENDENCE OF THE NEAR UV ABSORPTION CROSS SECTION OF HNO_{3}

\underline{J. A. PHILLIPS}, UCAR / NOAA Postdoctoral Research Fellow in Climate and Global Change, Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO 80309; V. VAIDA, Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO 80309.

In the atmosphere, HNO_{3} is an important reservoir species for the odd nitrogen (NO_{x}) and odd hydrogen (HO_{x}) families, both of which contribute to the destruction of stratospheric ozone. In the sunlit atmosphere, NO_{2} and OH are liberated by photolysis of HNO_{3}. Therefore, in order to model the distribution of these species, accurate values for the photodissociation cross section of HNO_{3} are necessary, particularly for wavelengths beyond 290 nm, which dominate the available light in the lower stratosphere. Previous investigations^{a} show significant discrepancies at wavelengths greater than 310 nm, and have also noted a marked temperature dependence in this region. In an effort to resolve these discrepancies, and determine if clusters of HNO_{3} are affecting the previous measurements, we have reexamined the near UV absorbance cross section from 240 to 350 nm at a wide range of pressures and temperatures. Implications of these results for the observed diurnal variation of [OH] in the lower stratosphere will be discussed.

aJ. B. Burkholder, R. K. Talukdar, A. R. Ravishankara, and S. Solomon, J. Geophys. Res. 98, 22937, (1993); O. Rattigan, E. R. Lutman, R. L. Jones, and R. A. Cox, Ber. Bunsenges. Phys. Chem. 96, 399, (1992); and references therein.